3. Cournot merger analysis. Suppose the low-cost firm wants to purchase the high-cost firm to become a monopoly with a marginal cost equal to 50. The Department of Justice (DOJ), however, is skeptical of this merger and fears that the monopoly will harm consumers. The low-cost firm asserts, however, that because it will be operating all production at a low cost, its price will be low and consumers will benefit. The DOJ hires you as a consultant to forecast the effect of the proposed merger. If the merger is approved, what price and quantity will be set by the new monopoly? Will consumers be helped or harmed by the merger? If the DOJ cares only about consumer surplus, should the merger be approved?
Answer #1

MR- MC MR- 800-80, ince ad Qove P-800-40 80 800-80 50 450 M9.5 New Bice P- 800- 408004 x8-45-425$ An 80 New P,B= (435, q345) C-S, x [800-486]x9345 = 1+548-135串 NOW then P= 31o, G-199.5 Jus 8 DoT cases about owy Gs, Meng e 3hould not be araved, dlhus Consumoa ahs not beuekited tcom Mene, buoz totat